# Adaptively Optimizing the Performance of HPX’s Parallel Algorithms

### v1

@@ Abstract
C++ Executors simplify the development of parallel algorithms by abstracting concurrency management across hardware architectures. They are designed to facilitate portability and uniformity of user-facing interfaces; however, in some cases they may lead to performance inefficiencies due to suboptimal resource allocation for a particular workload or not leveraging certain hardware-specific capabilities.
To mitigate these inefficiencies, we have developed a strategy based on cores and chunking (workload), and integrated it into HPX’s executor API. This strategy dynamically optimizes for workload distribution and resource allocation based on runtime metrics and overheads. In this paper, we introduce the model behind this strategy and evaluate its efficiency by testing its implementation (as an HPX executor) on both compute-bound and memory-bound workloads. The results show speedups across all tests, configurations, and workloads studied, offering improved performance through a familiar and user-friendly C++ executor API.
Keywords:
Asynchronous Many-Task (AMT) HPX Executors Parallel Algorithms.
@@ 1 Introduction
Managing parallelism and concurrency has become increasingly complex in software due to the heterogeneity of modern architectures. HPX provides executors to offer uniform APIs for execution of tasks across different architectures. While this abstraction eases development, it delegates many runtime metrics, such as number of cores or chunk sizes, to the execution policy, where poor choices can lead to performance degradation. For instance, even a straightforward embarrassingly parallel operation on a multi-core processor can be sensitive to the resource allocation strategy. While distributing work across all cores may be effective for larger input data, utilizing more cores for smaller input sizes may lead to performance degradation.
@@
Beyond that, there is a wide range of parallel algorithm types, including map-type algorithms (e.g. copy, fill, stencil updates, etc.), map-reduce-type algorithms (e.g. min_element, all_of, count, and prefix sums, etc.), to name a few. The performance of these algorithms can vary depending on several factors, such as the nature of the algorithm (whether it is compute-bound or memory-bound), the number of cores and threads, data partitioning, and the specific architecture (cache sizes, memory bandwidth, etc.). The challenge is that no single set of parameters can guarantee optimal performance across all scenarios. The optimal configuration for parallel execution is often highly dependent on the specific workload and runtime environment. In the context of parallelization, there are four factors to consider: Starvation:, Latencies:, Overheads: and, Waiting for contention resolution [1]
@@
This paper addresses these challenges by proposing a dynamic optimization approach integrated into HPX’s executor API. We propose an aggregated model which combines all the effects such as architectural specifics to get the best possible performance regarding the number of cores and chunk sizes. Our focus is on optimizing workload distribution and resource allocation through runtime adaptivity. This method aims to reduce overheads, such as those associated with parallelism management, while maintaining the simplicity of user-friendly APIs. We investigate the impact of different core counts and chunk sizes on two algorithms: adjacent-difference and artificial-work.
@@ 2 Motivation
Scientific simulation codes are often based on Cauchy problems, i.e., codes described by an initial data on an N-dimensional spatial grid, and accompanied by a system of differential equations used to iteratively evolve the data through time. These codes, common in science and industry, typically involve the use of finite differences to simulate spatial derivatives, operations that are very much like the adjacent-difference algorithm which calculates the difference between each pair of elements within a sequence. Such algorithms are inherently parallel and can benefit from parallelization of loops or regions of independent tasks. To parallelize loops, the computational grid is split into chunks (chunking), such that each chunk can be independently updated on a separate core. Determining the size of the chunks and a reasonable number of processing units to obtain the best possible execution times is—more often than not—left to the programmer. These metrics are application-specific and require measurement and tuning for each of the parallel constructs involved. Common off-the-shelf runtimes like OpenMP and the C++17 Standard library expect the user to tune and provide such metrics. The effectiveness of a parallelization strategy also depends on system and implementation-specific characteristics, e.g., the overhead per scheduled task, the amount of (possibly nested) parallelism intrinsic to the application, the underlying computing architecture, and many more factors [3].
To address this loop-tuning problem, we will develop a methodology for adaptively determining the optimal chunk sizes and the number of processing units and implement it for HPX’s parallel algorithms.
@@ 3 Theory
What is the ideal number of cores to use for a given workload?
To start, we define $T_{1}$ as the total time taken to execute a loop on a single thread without parallelism, and $T_{N}$ as the total time with $N$ number of threads (assuming $N>1$). We now assume that the loop can be perfectly parallelized apart from a constant overhead, $T_{0}$.
@@
Alternatively, we can derive the relation between speedup (which measures how much faster a parallel algorithm runs in compare to a sequential one), efficiency (which measures how well a system utilizes parallelism), and number of processors to calculate the optimal number of processors from first principles.
We can now compute speedup as follows:
@@
Thus far, the speedup formula is generic to any model of computation, including Amdahl’s Law[4] and Gustafson’s Law[5]. However, substituting Equation 1 into Equation 2, we obtain the following:
@@
So the "Overhead Law" is differs from Amdahl’s Law by a constant factor. We point out that, unlike the other laws, this equation is not valid for the case $N=1$ since Equation 1 only applies when $N>1$.
@@
Using Equation 1 and substituting $T_{N}$ based on $T_{1}$, simplifying them, we can solve it for N:
@@
We will choose an efficiency (E) of 95% as before.
@@
In summary, this approach allows us to calculate $N_{C}$, the number of cores that we can make use of without exceeding our parallel efficiency goal.
Once the optimal core count for a given workload is determined, we need to find the optimal chunk size. To find that, We apply the following formula to determine the ideal amount of work per core:
$N_{C}\>=\>{\frac{T_{1}}{T_{opt}}}$ (8)
where as before, $T_{1}$ is the total time taken per workload and $T_{opt}$ is the time for minimum work per core. If we find the amount of work where parallelization with 2 cores gives a speedup of $1.9$ (a $95\%$ efficiency), based on our experiments we can, with the given formula, calculate $T_{opt}$. This optimal time represents the amount of work per core for $95\%$ efficiency. we derive the following formula using Eq. 8 to calculate the minimum workload for each chunk:
$T_{m}\>=\>{\frac{T_{1}}{N_{C}*C}}$ (9)
Where $N_{C}$ is the optimal number of cores calculated in Eq. 8, $C$ is chunks-per-core (which is equal to 8 based on the experiments), $T_{1}$ is the overall time per workload and $T_{m}$ is the amount of work per chunk. Assuming that $T_{1}$ and $T_{m}$ are both proportional to the number of elements they contain, we can derive an equation for the chunk size ($N_{CH}$). The value of $N_{E}$ is the number of elements in our workload.
$N_{CH}\>=\>{\frac{N_{E}}{N_{C}*C}}$ (10)
This equation ensures that $C=8$ chunks per core are used for any workload, with the chunk size always being at least $T_{m}$. In section 4 we will show how this time parameter is utilized to design the chunk size.
@@ 4 Implementation
Now that we know how to compute the ideal number of cores, $N_{C}$, we turn to the implementation. We use HPX because it is open source, easily customizable, and well-optimized. For this work, we will consider the parallelization of algorithms and loops, scheduling work in a manner similar to the familiar static scheduling of OpenMP.
@@
4.1 HPX Algorithms and Customization
HPX is a C++ Standard Library for parallelism and concurrency [6, 7]. HPX is implemented as a lightweight user-level task manager running on top of kernel threads [8].
Execution policies are objects that can be provided as the first argument to the standard algorithms of C++. The C++17 and C++20 Standards introduce four execution policy types [9, 10].
  • std::execution::seq requires that a parallel algorithm’s execution not be parallelized.
  • std::execution::par indicates that a parallel algorithm’s execution may be parallelized. Within this policy, all the different scheduling types
    from OpenMP are available through objects called execution parameters.
@@ 4.1 HPX Algorithms and Customization
  • std::execution::par_unseq indicates that a parallel algorithm’s execution may be parallelized, vectorized, or migrated across threads.
@@
In HPX, we have implemented all of the above with correct semantics in the namespace hpx::execution. However, the community quickly realized that the execution policies alone do not provide sufficient flexibility for controlling the (possibly parallel) execution environment. Early discussions in the ISO standardization committee documented the idea of standardizing ‘unified executors’, which was later abandoned [11]. In HPX, we have used the ideas outlined in that paper to develop the extensive customization mechanisms that we describe below.
@@
All of the parallel algorithms in HPX rely on internally invoking a number of customization points that can be overloaded by either the executor or the execution parameters object. We call a customization point a function that can be overloaded by an external user-defined function or object such that it is selected at compile time as a replacement for an internal, predefined default implementation. Customization point objects are usually function object instances that fulfill the two objectives of a) unconditionally trigger (conceptified) type requirements on the arguments of that function, and b) dispatch to the correct function via argument dependent lookup (ADL). In C++, these are often a library feature that adds concept checking resulting in, e.g. clearer compilation error messages in case of erroneous template instantiations. We describe the set of customization points used by HPX’s standard algorithm implementation relevant to this work in Section 4.2. In HPX, we rely on the tag_invoke methodology [12] for the implementation of our customization points.
4.2 Customization Points for HPX’s Parallel Algorithms

### v2

# A New Execution Model and Executor for Adaptively Optimizing the Performance of Parallel Algorithms Using HPX Runtime System

@@ Abstract
Developing parallel algorithms efficiently requires careful management of concurrency across diverse hardware architectures. C++ executors provide a standardized interface that simplifies the development process, allowing developers to write portable and uniform code. However, in some cases, they may not fully leverage hardware capabilities or optimally allocate resources for specific workloads, leading to potential performance inefficiencies. Building on our earlier conference paper [Adaptively Optimizing the Performance of HPX’s Parallel Algorithms], which introduced a preliminary strategy based on cores and chunking (workload), and integrated it into HPX’s executor API, that dynamically optimizes for workload distribution and resource allocation, based on runtime metrics and overheads, this paper, introduces a more detailed model of that strategy. It evaluates the efficiency of this implementation (as an HPX executor) across a wide range of compute-bound and memory-bound workloads on different architectures and with different algorithms. The results show consistent speedups across all tests, configurations, and workloads studied, offering improved performance through a familiar and user-friendly C++ executor API. Additionally, the paper highlights how runtime-driven executor adaptations can simplify performance optimization without increasing the complexity of algorithm development
@@ 1 Introduction
Managing parallelism and concurrency has become increasingly complex in software due to the heterogeneity of modern hardware. HPX provides executors to offer uniform APIs for execution of tasks across different architectures. While this abstraction eases development, it delegates many runtime metrics, such as number of cores or chunk sizes, to the execution policy, where poor choices can lead to performance degradation. For instance, even a straightforward embarrassingly parallel operation on a multi-core processor can be sensitive to the resource allocation strategy. While distributing work across all cores may be effective for larger input data, utilizing more cores for smaller input sizes may lead to performance degradation.
@@
Beyond hardware considerations, there is a wide range of parallel algorithm types, including map-type algorithms (e.g. copy, fill, stencil updates, etc.), map-reduce-type algorithms (e.g. min_element, all_of, count, and prefix sums, etc.), to name a few. The performance of these algorithms can vary depending on several factors, such as the nature of the algorithm (whether it is compute-bound or memory-bound), the number of iterations, the number of cores and threads, data partitioning, and the specific architecture (cache sizes, memory bandwidth, etc.). The challenge is that no single set of parameters can guarantee optimal performance across all scenarios. The optimal configuration for parallel execution is often highly dependent on the specific workload and runtime environment.
Ideally, parallel speedup should scale linearly with the number of processors [1]. However, as Amdahl’s Law [2] has shown, the maximum speedup of a process on multiple processors is limited by its sequential portion. This limitation can be mitigated by increasing the amount of parallel work as more processors are available, as described by Gustafson’s Law [3]. More generally, in the context of parallelization, there are four factors to consider:
  • Starvation: Occurs when there’s not enough concurrent work available, leading to reduced resource utilization.
  • Latencies: Refers to time-distance delays in accessing remote resources.
  • Overheads: Refers the work required for managing parallel operation, a feature absent in sequential codes.
  • Waiting for contention resolution: Delays due to the unavailability of shared resources.
All these factors contribute to reduced parallel efficiency. This paper focuses on reducing overheads. We study the effects on performance of workload distribution (chunk size) and the number of processors utilized.
@@
This paper addresses these challenges by proposing a dynamic optimization approach integrated into HPX’s executor API. We propose an aggregated model that combines effects such as architectural specifics to get the best possible performance regarding the number of cores and chunk sizes. Our focus is on optimizing workload distribution and resource allocation through runtime adaptivity. Unlike approaches that try to break down overhead into many separate factors (cache effects, memory bandwidth, etc.), our model combines all those contributing factors as a single, combined overhead value $T_{0}$. We measure this combined overhead per loop invocation on each target architecture instead of attempting to calculate it. This method aims to reduce overheads, such as those associated with parallelism management, while maintaining the simplicity of user-friendly APIs. In our previous paper [4], we introduced an adaptive performance model and examined the impact of varying core counts and chunk sizes on two algorithms: adjacent-difference and artificial-work. In this paper, we extend our analysis by running additional benchmarks on diverse architectures to validate the model’s architecture‑independence. We also explore a wider range of chunking configurations to confirm that our identified optimal chunking remains effective. Finally, we evaluate our enhanced approach on an algorithm implemented in ChplX [5], which is a source to source translator that converts a Chapel [6] program into a C++ program using HPX.
@@ 2 Motivation
Scientific simulation codes are often based on Cauchy problems, i.e., codes described by an initial data on an N-dimensional spatial grid, and accompanied by a system of differential equations used to iteratively evolve the data through time. These codes, common in science and industry, typically involve the use of finite differences to simulate spatial derivatives, operations that are very much like the adjacent-difference algorithm that calculates the difference between each pair of elements within a sequence. Such time integration methods and algorithms often expose sufficient parallel work that can benefit from parallel computational constructs like parallel loops or parallel regions of independent tasks. On CPUs, parallelizing loops is performed by splitting (chunking) the computational grid such that the parts of the grid can be independently and concurrently scheduled on separate processing units (cores). Determining the size of those chunks and the number of processing units to utilize that result in the best possible execution times is—more often than not—left to the programmer. Selecting the optimal parameters for these parallel loops requires application specific measurements and manual tuning for each of the parallel constructs involved. Common off-the-shelf runtimes like OpenMP and the C++17 Standard library expect the user to tune and provide such metrics. This problem is aggravated when using library solutions like the C++17 Standard algorithms, as the optimal size of the chunks and optimal number of cores depend on the size of the computational space and the amount of work in the user-supplied element access functions passed to those algorithms. Naturally, this amount of work is unknown to the implementer of the parallel algorithms. The optimal values for the parallelization parameters also depends on system- and implementation specific characteristics, e.g. the overhead per scheduled task, the general amount of—possibly nested—parallelism intrinsic to the application, the underlying computing architecture, and many more factors [8].
To address this problem, we describe a methodology for adaptively determining the optimal chunk sizes and the number of processing units used for HPX’s parallel C++ Standard algorithms. We have developed a mathematical model that we use to dynamically find the number of cores and chunk sizes for map-type parallel algorithms that result in the best possible execution performance of that algorithm. We have implemented this model as an HPX execution policy and studied its performance on a variety of algorithms.
In order to achieve this goal, we measure the time to execute the first invocation of the loop. The measurement is passed to the model to calculate the best core-chunk configuration. The result is used by all subsequent invocations of the same parallel loop. This approach reduces the measurement overhead required to find the time per loop execution for a particular platform and execution environment by amortizing this overhead over all subsequent calls of the loop.
The required functionality is encapsulated in a simple C++ type that serves as the customization point for HPX’s parallel algorithms, and so it does not require us to modify the algorithms’ implementation. This will be explored in greater depth in section 4.1.1.
Some recent studies also address chunking, scheduling, and load balancing for parallel loops. For example, the Distributed Chunk Calculation Approach (DCA) compares centralized vs distributed chunk size calculation under CPU slowdown scenarios, showing gains when chunk sizing is made more adaptive to runtime conditions [9]. Likewise, Chunks and Tasks provides abstractions for splitting both data and work into chunks and tasks with dynamic work distribution, which helps with irregular workloads and balancing [10]. In a similar work  [11] introduce the Heartbeat Compiler (HBC), which translates C/C++ programs with nested parallelism into binaries capable of automatic granularity control, effectively managing overheads in irregular workloads
While these approaches enhance parallelism through adaptive chunking and scheduling, they do not incorporate runtime measurements to derive overhead and sequential execution times. Our work differs by measuring early loop runtimes to estimate these parameters, enabling dynamic adaptation of both core count and chunk size based on empirical data, thereby optimizing performance in real-time.
@@ 3 Theory
To start, we define $T_{1}$ as the total time taken to execute a loop on a single thread without parallelism, and $T_{N}$ as the total time with $N$ being the number of threads (assuming $N>1$) used to execute the loop. We now assume that the loop can be perfectly parallelized apart from a constant overhead, $T_{0}$.
@@
Alternatively, we can derive the relation between speedup (that measures how much faster a parallel algorithm runs compared to a sequential one), efficiency (that measures how well a system utilizes parallelism), and number of processors to calculate the optimal number of processors from first principles.
We can now compute the speedup as follows:
@@
Thus far, the speedup formula is generic to any model of computation, including Amdahl’s Law and Gustafson’s Law. However, substituting Equation 1 into Equation 2, we obtain the following:
@@
So the "Overhead Law" differs from Amdahl’s Law by a constant factor. We point out that, unlike the other laws, this equation is not valid for the case $N=1$ since Equation 1 only applies when $N>1$.
@@
After using Equation 1 and substituting for $T_{N}$, we can simplify and solve for N:
@@
We will, somewhat arbitrarily, choose an efficiency (E) of 95%. A different choice could be made, and the remainder of the analysis would change accordingly.
@@
Alternatively, we can compute the optimal core count, $N$, using $T_{opt}$.
$N\>=\>{\frac{T_{1}}{T_{opt}}}$ (8)
where as before, $T_{1}$ is the total time taken per workload and $T_{opt}$ is the time for minimum work per core. If we find the amount of work where parallelization with 2 cores gives a speedup of $1.9$ (a $95\%$ efficiency), based on our experiments we can, with the given formula, calculate $T_{opt}$. This will result in $T_{opt}=19T_{0}$ and provide the same answer as Equation 7.
Once the optimal core count for a given workload is determined, we need to find the optimal chunk size. To find that, we use the following formula to calculate the minimum workload for each chunk:
$T_{m}\>=\>{\frac{T_{1}}{N*C}}$ (9)
Where $C$ is chunks-per-core (which is equal to 8 based on the experiments), $T_{1}$ is the overall time per workload and $T_{m}$ is the amount of work per chunk. Assuming that $T_{1}$ and $T_{m}$ are both proportional to the number of elements they contain, we can derive an equation for the chunk size ($N_{Chunk}$). The value of $N_{E}$ is the number of elements in our workload.
$N_{Chunk}\>=\>{\frac{N_{E}}{N*C}}$ (10)
This equation ensures that $C=8$ chunks per core are used for any workload, with the chunk size always being at least $N_{Chunk}$.
The remaining question is ’How can $T_{0}$ and $T_{1}$ be estimated from measurements? To obtain concrete values for $T_{0}$ (the overhead) and $T_{1}$ (time for sequential work), we measure $T_{N}$ for several thread counts N>1. We assume the model
$T_{N}=T_{0}+T_{1}/N$ (11)
and use ordinary least squares regression to find $T_{0}$ and $T_{1}$ that best fit the observed data (minimizing the sum of squared differences between predicted and measured runtimes) [12]. Once these parameters are estimated, they are substituted into Equations 1 to Equation 10 to drive predictions of speedup, efficiency, and optimal configurations in our model.
Another practical way to measure $T_{0}$ is to measure the loop once in sequential mode and once in parallel with one core. Then the difference between these two values is the $T_{0}$ per core which we need for our model. This is identical to the approach we mentioned above.
@@ 4 Implementation
Now that we know how to compute the ideal number of cores, $N$, we turn to the implementation. We use HPX because it is available as an open source library, easily customizable, and well-optimized. For this work, we will consider the parallelization of algorithms and loops, scheduling work in a manner similar to the familiar static scheduling of OpenMP.
@@
4.1 HPX
HPX is a C++ Standard Library for parallelism and concurrency [13, 14]. HPX is an asynchronous many-task runtime (AMT) system that is built on the ParalleX execution model [15]. It is implemented as a lightweight user-level task manager running on top of kernel threads [16]. It is widely known that thread creation and destruction managed by the operating system are expensive and reserve lots of memory [16]. For that reason, HPX creates and binds one kernel thread (running a user-level queue with work-steeling) to each core [15]. User-level threads (i.e. tasks) can be shared among these queues and executed in parallel. HPX thereby abstracts the writing of parallel code from the number of cores available on the machine.
In addition, HPX suspends threads instead of blocking for mutexes, waiting on futures, etc. This means that HPX is able to assign one worker per core and keep the workers busy even for applications that, semantically, are blocking.
The HPX asynchronous programming model exposes a C++ standard API entirely conforming to interfaces as defined by the C++ Standard and extends the standardized APIs by providing distributed and heterogeneous computing based on Futures, which makes HPX portable and uniformly usable for local and remote parallelism [16].
In this paper we focus on HPX’s customization points for the control of parallelism within the algorithms to implement our newer, more dynamic method.
4.1.1 HPX Algorithms and Customization
Execution policies are objects that can be provided as the first argument to the standard algorithms of C++. The C++17 and C++20 Standards introduce four execution policy types [17, 18]. All of these are used for unique predefined type instances to disambiguate parallel algorithm overloading:
  • std::execution::seq requires that a parallel algorithm’s execution not be parallelized. The invocations of element access functions in parallel algorithms are indeterminately sequenced in the calling thread.
  • std::execution::par indicates that a parallel algorithm’s execution may be parallelized. The invocations of element access functions are permitted to execute in either the invoking thread or in a thread implicitly created by the library and are indeterminately sequenced with respect to each other.
  • std::execution::par_unseq indicates that a parallel algorithm’s execution may be parallelized, vectorized, or migrated across threads. The invocations of element access functions are permitted to execute in an unordered fashion in unspecified threads, and unsequenced with respect to one another within each thread.
@@ 4.1.1 HPX Algorithms and Customization
In HPX, we have implemented all of the above with correct semantics in the namespace hpx::execution. However, while implementing all of the parallel algorithms as specified by the C++ Standard, the community quickly realized that the execution policies alone do not provide sufficient flexibility for controlling the execution environment. Early discussions in the ISO standardization committee documented the idea of standardizing ‘unified executors’, which was later abandoned [19]. In HPX, we have used the ideas outlined in that paper to develop the extensive customization mechanisms that we describe below.
@@
All of the parallel algorithms in HPX rely on internally invoking a number of customization points that can be overloaded by either the executor or the execution parameters object (see Section 4.1.2). We call a customization point a function that can be overloaded by an external user-defined function or object such that it is selected at compile time as a replacement for an internal, predefined default implementation. Customization point objects are usually function object instances that fulfill the two objectives of a) unconditionally trigger (conceptified) type requirements on the arguments of that function, and b) dispatch to the correct function via argument dependent lookup (ADL). In C++, these are often dependent on a library feature that adds concept checking, resulting in, e.g. clearer compilation error messages in case of erroneous template instantiations. We describe the set of customization points used by HPX’s standard algorithm implementation relevant to this work in Section 4.1.3. In HPX, we rely on the tag_invoke methodology [20] for the implementation of our customization points.
4.1.2 HPX Executors and Execution Parameter Objects
In a previous paper [4], we briefly explained HPX and Customization points. In this paper, we will elaborate on these topics and provide more details about executors and parameter objects which are crucial to our implementation.
Executors provide a uniform interface for work creation and task scheduling by abstracting underlying resources where work physically executes. Underlying resources could be, for instance, a thread pool, SIMD units, GPU runtimes, or simply the current thread. In general, we call such resources execution contexts. As lightweight handles, executors impose uniform access to execution contexts. The uniformity of the API exposed by executors enables control over where work executes, even when it is executed and scheduled indirectly behind library interfaces. All of HPX’s APIs related to executing threads, in particular all parallel algorithms, rely on executors for their thread scheduling needs. In a nutshell, executors are used to describe when, where, and how to execute a computation.
One notable advantage of utilizing executors is the ability to maintain optimizations external to and independently of the algorithm implementations, empowering users to decide when and where to leverage these optimizations. Consequently, the algorithm implementations themselves remains untouched.
In HPX, several executors have been implemented to control various functionalities, such as restricting execution to a subset of the available processing units (cores) or to encapsulate different thread-scheduling implementations. The focus of this paper centers on an executor responsible for managing an optimal number of cores and optimal chunk sizes used while parallelizing the execution of map-style and map-reduce-style parallel algorithms.
Execution Parameter Objects provide a uniform way of customizing additional execution parameters, such as chunk sizes, priorities, thread affinities, or the number of processing units to be used for running a parallel algorithm. In this paper, we present an implementation of a new executor parameters object that adaptively controls the chunk sizes and the number of cores used to execute a given parallel algorithm.
The HPX execution policies expose additional APIs that allow the programmer to attach an executor (.on()) and/or execution parameter objects (.with()) as shown in Listing 1. Both APIs create new execution policies that encapsulate the given executor and/or execution parameters object. This design decision enabled a seamless extension of the well-documented parallel algorithm interface while providing deep and fine-grain customization capabilities to the user.
// use default execution policy execution::par
// relying on default executor
std::vector<double> d = { ... };
fill(execution::par, begin(d), end(d), 0.0);
// rebind par to an user-defined executor
// using .on()
user_defined_executor ex = ...;
fill(execution::par.on(ex), ...);
// rebind par to an user-defined executor and
// a user-defined execution parameters object
// using .on() and .with()
user_defined_params param = ...
fill(execution::par.on(ex).with(param), ...);
Listing 1: Customization examples for HPX execution policies
In this paper, we introduce an executor that optimizes the number of cores and chunks for various workloads. Users can easily integrate this executor into their parallel algorithms using .with().
Our experiments focus on the parallel implementation of the adjacent_difference algorithm in HPX. Listing 2 demonstrates how users can attach this executor, referred to as acc, to the algorithm. Note that we pass acc explicitly by reference to support the use case described in Section 2. The first invocation of the algorithm will measure the time per loop execution and store the result in acc, thus making the time available to all subsequent invocations of the algorithm.
adaptive_core_chunk_size acc;
adjacent_difference(
execution::par.with(std::ref(acc))),
arr.begin(), arr.end(), res1.begin());
Listing 2: Integrating the new execution parameters object (acc) with the adjacent-difference algorithm in HPX
4.1.3 Customization Points for HPX’s Parallel Algorithms